# ImageNot: A contrast with ImageNet preserves model rankings

### v1

@@ 3 Constructing ImageNot
The second significant not is the accuracy of image-synset pairs. In ImageNet, an average precision of 99.7% is achieved on a total of 80 synsets randomly sampled at different depths of the WordNet tree. Candidate images were collected from the internet by querying image search engines, and then human annotators were employed to verify each candidate image for a given synset, yielding relatively accurate image-synset pairs. In contrast, we deliberately make no such cleaning steps in ImageNot. We rely entirely on similarities between embeddings from CLIP (image and caption similarities) and RoBERTa (caption and class definition similarities) to select class and image pairs. Consequently, ImageNot’s labels are far more noisy than ImageNet, which utilizes human annotators. This noisiness is noticeable in the accuracy observed when training on ImageNot – the best model achieves around 50%, which is drastically better than chance (0.1%), but lower than what the best model achieves on ImageNet (around 85%). We also see the effect of this process in the distribution of inter-class image similarities between ImageNot and ImageNet illustrated in Figure 6; the similarity in images corresponding to an ImageNet class tends to be higher than those in an ImageNot class.
Figure 7: Number of synset siblings.
Similarities in ImageNot and ImageNet WordNet subtrees.
@@
Despite the differences in ImageNot, the subset of the WordNet trees spanned by their synsets are structurally similar. They share approximately the same number of nodes and edges, have similar depth, density and modularity, and share a 23% overlap in the parents of the datasets’ synsets – illustrated in Figure 7 (more details about the WordNet subtrees can be found in Appendix B.1 Table 3).
@@ 4 ImageNot preserves model rankings
(a) Models initialized with random weights. Train models on a dataset. Evaluate models on the same dataset’s held-out test set.
(b) Relative progress of each model architecture. For a set of test accuracies $X$, relative progress is $X_{i}/X_{\text{AlexNet}}$.
(c) Test accuracy compared to ImageNet test accuracy for each model architecture.
Figure 8: ImageNet vs. ImageNot. We also show 95% confidence intervals for the regression estimates. Robust corresponds to average accuracy across many test-time distribution shifts.
Our experimental design is centered around evaluating whether ImageNot maintains the model rankings established by ImageNet. Specifically, we want to know: if a new architectural advancement was evaluated on ImageNot instead of ImageNet, would it still be recognized as a significant development? Our focus is on a collection of model architectures that represent significant milestones in the evolution of ImageNet models. This includes AlexNet (Krizhevsky et al., 2012), VGG (Simonyan and Zisserman, 2014), ResNet (He et al., 2016), DenseNet (Huang et al., 2017a), EfficientNet (Tan and Le, 2019), and ConvNeXt (Liu et al., 2022). We motivate this set in section 2. For each model, we select the configuration that achieved the highest reported performance on ImageNet, such as ResNet152 over ResNet18. These models are then trained and evaluated on ImageNot, allowing us to directly compare their rankings with those obtained on ImageNet.
@@
We also investigate model rankings on the CIFAR10 dataset, which consists of 60,000 32x32 colored images of 10 classes, with 5000 train images per class and 1000 test images per class (Krizhevsky et al., ; Krizhevsky et al., 2009). The learning task is 10-class image classification – Airplane, Automobile, Bird, Cat, Deer, Dog, Frog, Horse, Ship, and Truck. We note that ImageNet contains 50 test images per class (Deng et al., 2009) while ImageNot contains 100.
We consider three types of rankings: (i) Standard. We initialize each model with random weight, then train each model on a dataset, and finally, we evaluate each model on the same dataset’s test set. (ii) Fine-Tuning. We initialize each model with pre-trained on ImageN[o/e]t weights, then we fine-tune the model end-to-end (all weights updated) on CIFAR10, and finally, we evaluate it on CIFAR10’s test set. (iii) Transfer Learning. We initialize each model with pre-trained ImageN[o/e]t weights, then we update only the model’s last fully connected layer while training on CIFAR10, and finally, we evaluate it on CIFAR10’s test set.
@@ 4.1 Model development
ImageNet. Our reference rankings are derived from pre-trained models available on Torch Hub (Marcel and Rodriguez, 2010; Paszke et al., 2019). These models establish a ranking, which we consider our factual observations. In decreasing order of performance, the ranking is as follows: EfficientNet V2 Large, ConvNeXt Large, ResNet 152, DenseNet 161, VGG 19, ResNet 18, and AlexNet. For additional specifics regarding these pre-trained models, please refer to Appendix A.1.
Training.
We also utilize the implementation and pre-trained model weights from torch hub (Marcel and Rodriguez, 2010; Paszke et al., 2019), when applicable. For the first ranking category (no pretraining), we initialize each model with randomized weights and train end-to-end, meaning all model weights are updated with each backward pass.
On avoiding self-fulfilling prophecies.
There is a subtle pitfall that one may fall into if not careful – the inadvertent confounding effect of knowing the desired ImageNet ranking. Specifically, an experimenter aware of the expected ranking might unconsciously adjust the process until a particular model is in the right place. This iterative tuning could falsely lead to a preservation of the known rankings. To mitigate this risk, we implemented several a priori restrictions on the experimenter’s approach for each ranking modality:
  1. 1.
    The experimenter can only tune the learning rate, learning rate scheduler hyper-parameters, and weight decay constant.
  2. 2.
    There is a shared hyper-parameter set across all models.
  3. 3.
    There is a fixed # of training runs that can be launched for the same maximum epoch.
  4. 4.
    The fixed hyper-parameter set can only be changed (expanded) if some model(s) do not train.
See further discussion in Appendix A.2. This cautious approach introduces the primary limitation of our work. To avoid over-tuning and maintain objectivity, we refrain from employing extensive model tuning or leveraging all modern performance-enhancing techniques, such as data augmentation, label smoothing, etc. As a result, it is possible that we do not end up with the optimal training configuration for each model (Figure 8(c)).
@@ 4.2 Model Rankings
ImageNot preserves model rankings and relative improvements.
@@
Figure 8(a) illustrates that the model rankings between ImageNot and ImageNet are remarkably consistent. In Figure 8(b), we compare each model’s performance improvement relative to AlexNet, the model that kick-started the deep learning era on ImageNet in 2012. This comparison reveals a notable parallel in the progression of accuracy improvements across both ImageNet and ImageNot. Not only do the ordinal rankings of models remain unchanged, the degree of performance enhancement is also consistent. In simpler terms, this suggests that if a new architecture like EfficientNet (Tan and Le, 2019) had been introduced in a different universe focused on ImageNot instead of ImageNet, we would have witnessed a comparable impressive advancement. The external validity of ImageNet improvements holds! The ImageNet vs. ImageNot test accuracies have a linear fit with a slope of 0.46 and an adjusted $R^{2}$ of 0.98. ImageNet vs. CIFAR10 test accuracies have a linear fit with a slope of 0.18 and an adjusted $R^{2}$ of 0.93. We note that these slopes should not be interpreted the same way as those in Recht et al. (2019) because we compare test accuracies across datasets, i.e., different ranges in accuracy across models.
(a) ImageN[o/e] pre-trained models. Train on CIFAR10. Evaluate on the CIFAR10. Similar to previous rankings, changes occur between architectures with overlapping error bars.
@@
(b) Fine-tuned on CIFAR10 test accuracy compared to standard ImageNet test accuracy for each model architecture.
@@
(c) Transfer Learning (fine-tune only last fully-connected layers) on CIFAR10 test accuracy compared to standard ImageNet test accuracy for each model architecture.
@@
Figure 9: ImageNet vs. ImageNot. Fine-tuning and Transfer Learning. Robust corresponds to average accuracy across many test-time distribution shifts. We also show 95% confidence intervals for the regression estimates.
@@
Figure 8(c) highlights the limitations of relying solely on the external validity of accuracy as the primary metric for the effectiveness of a benchmark. One may look at the drop in accuracy from ImageNet to ImageNot and conclude that ImageNet results have no external validity but this would be a rather narrow and inaccurate conclusion. From the standpoint of benchmarking, the more relevant observation and notion of external validity is that ImageNet accuracy is a strong predictor of ImageNot accuracy (Figure 8(c)), despite the significant difference between the two datasets.
@@
For each CIFAR10 ranking, we do three random restart restarts and compute the standard errors for each model’s test accuracy. Given that we opted to include architectures that achieve very similar accuracies on the ImageNet test set (such as EfficientNet-V2-L/ConvNeXt-L or ResNet152/DenseNet161), we expect potential training variance to affect the model, particularly given the size of CIFAR10. For instance, when training CIFAR10 from randomly initialized weights, EfficientNet-V2-L shows a slightly lower average test accuracy than ConvNeXt-L across the random restarts. However, their error bounds intersect, so the ranking is preserved up error margins.
@@
ImageNot improvements yield improvements in fine-tuning and transfer learning.
When fine-tuning pre-trained ImageN[o/e]t models on CIFAR10, we observe that model rankings are still preserved (Figure 9(a)). Within a fixed hyperparameter set, however, we find that some of the models, particularly those with the highest accuracies, have high variance. Nevertheless, both the upper bounds and lower bounds of the error margins increase as the accuracy on ImageNet increases (Figure 9(b)).
Furthermore, despite the arbitrary nature of the ImageNot dataset, we observe that the representations it induces are still transferable – the accuracy of a model architecture on ImageNot is still a good predictor of the accuracy achieved by retraining the last layer of a model pre-trained on ImageNot on CIFAR10, (Figure 9(c)). However, we note that the ImageNot features are notably more noisy than ImageNet features when fine-tuning. This suggests that the noisiness of ImageNot does have some negative downstream effects on transfer learning. Nevertheless, we generally observe a positive trend in ImageNot test accuracy and transferability to CIFAR10.
Rankings are preserved under robustness.
We also find that model rankings remain consistent when we look at average performance across 17 distribution shifts applied to the ImageN[o/e]t test set, e.g., random blurring, cropping, and rotation. In this setting, our findings remain the same on average across these distribution shifts but with a consistent decrease in test accuracy for each model. This further shows that metrics can be a misleading definition of external validity as we observe changes in accuracy across different shifts. More details on rankings under robustness are in Appendix B.2.
Safety and ethical limitations of web-crawled datasets.
While our results suggest that web-crawled datasets with minimal human intervention may be sufficient for benchmarking progress in architecture development, other ethical considerations are a concern. As illustrated by Thiel (2023), the internet contains many illegal images and images of illegal activity. Particularly, Thiel (2023) find evidence of Child Sexual Abuse Material (CSAM) in LAION-5B. We guard against this by using a NSFW language classifier (Song et al., 2021) to remove synsets that may be associated with unethical images on top of the default LAION-5B NSFW filter. We did not manually inspect every image in ImageNot, but we employed our model-based filtering and design choices to avoid problematic images as best as possible. Nevertheless, we do not release the images or URLs for ImageNot, but we provide ImageNot class names, a detailed description of our generation procedure, and our code so that our results can be reproduced.
@@ 5 Conclusion
Despite the significant differences in ImageNet and ImageNot, we find consistent preservation of model rankings across both datasets. Moreover, this consistency extends to the fine-tunability of pre-trained models and the transferability of pre-trained image embeddings. For benchmarking, we propose that providing robust model rankings should be a hallmark of a good benchmark. Furthermore, our results suggest that concerns about the external validity of model developments on ImageNet may be overstated.
Many domain experts attribute the rapid advancement in object recognition models to the ImageNet benchmark. From this work, two takeaways emerge to guide and motivate future benchmarks. First, a benchmark that produces robust model rankings, even if accuracy degrades under distribution shifts, should be considered effective. Second, the emphasis on clean labels and accuracy numbers may not be necessary to measure progress effectively, given that the noise is unbiased.
The exact extent to which a benchmark needs to align with real-world image distributions to measure progress effectively remains a question for future work. Our findings indicate that this alignment might need to be less stringent than currently perceived in the machine learning community.
Additional future work includes changing additional design choices of ImageN[o/e]t, e.g., moving away from wordnet noun synsets and using other attributes of or relationships between concepts in images. Another direction for future work is to examine the properties considered in this work for fine-tuning ImageNet and ImageNot-type models on other downstream tasks such as image segmentation.
@@ Appendix A Architectures and Training
AlexNet (Krizhevsky et al., 2012) demonstrated the efficacy of deep learning and many architectural advances such as convolutional filters, rectified linear units (ReLUs) as activation functions, and more. Additionally, its relative success on the ILSVRC 2012 is considered to have revitalized interest in neural networks, sparking a wave of research in deep learning. VGG (Simonyan and Zisserman, 2014) then implemented a simpler and deeper architecture with smaller convolutional filters, along with regularization techniques such as dropout, which yielded improved accuracy and generalization. ResNet (He et al., 2016) further improved upon the architecture of these deep learning models with the introduction of residual blocks, which allowed for the training of deeper networks while mitigating the vanishing gradient problem and achieving superior performance to its predecessor. DenseNet (Huang et al., 2017a) proposed to connect each layer to every other layer to maximize information flow between layers. EfficientNet (Tan and Le, 2019) then improved upon its predecessors with compound and systematic scaling of depth, width, and resolution, as well as combinations of novel optimization and learning strategies, without increasing computational cost. ConvNeXt (Liu et al., 2022) aimed to compete with the state-of-the-art performance of Vision Transformers (ViTs; (Dosovitskiy et al., 2020)) by combining components of preceeding ConvNet architectures, e.g., residual blocks and kernel sizes.
A.1 Torch Hub Pretrained Models
We use pretrained models from Torch Hub; we summarize their training details below (see PyTorch Team (2024) for additional details).
Table 1: Default Hyperparameters
Parameter Default Value
batch size 32
epochs 90
lr 0.1
momentum 0.9
weight decay 1e-4
lr step size 30
lr gamma 0.1
AlexNet, VGG19.
Learning rate: 1e-2. Learning rate scheduler: StepLR.
ResNet, DenseNet. Learning rate scheduler: StepLR.
EfficientNet V2 L. Learning rate: 5e-1. Learning rate scheduler: Cosine annealing with linear warmup. Weight decay: 2e-5. EfficientNet V2 L also utilizes data augmentation, label smoothing, and other regularization techniques.
ConvNeXt Large. Learning rate: 1e-3. Learning rate scheduler: Cosine annealing with linear warmup. ConvNeXt also utilizes data augmentation, label smoothing, and other regularization techniques.
For both finetuning and transfer learning from ImageNet, we use these pretrained weights as our starting point.
A.2 Trained Models
We note that for ImageNot, the extent of our parameter tuning is significantly less than these pretrained ImageNet models (Table 2). Despite this, model rankings are still preserved. In general, when we train CIFAR10, we use the same set of hyperparameters, given that we can perform a broader search. This includes training from randomly initialized weights, finetuning, and transfer learning.
For ImageNet, the tuning of these parameters we manual and not search-based – each new hyperparameter was added based on the loss curves of models that did not sufficiently train and domain expertise of which hyperparameters can be changed to allow for training.

### v2

@@ 3 Constructing ImageNot
The second significant not is the accuracy of image-synset pairs. In ImageNet, an average precision of 99.7% is achieved across 80 synsets randomly sampled at different depths of the WordNet tree (deng2009imagenet). Candidate images were collected from the internet by querying image search engines, and then human annotators were employed to verify each candidate image for a given synset, yielding relatively accurate image-synset pairs. In contrast, we deliberately make no such cleaning steps in ImageNot. We rely entirely on similarities between embeddings from CLIP (image and caption similarities) and RoBERTa (caption and class definition similarities) to select class and image pairs. Consequently, ImageNot’s labels are far noisier than ImageNet, which utilizes human annotators. We note that the ‘noise’ encompasses both incorrect labels due to automated curation and the greater natural variability of the images in ImageNot. This noisiness is noticeable in the accuracy observed when training on ImageNot – the best model achieves around 60%, which is drastically better than chance (0.1%) but lower than what the best model achieves on ImageNet (around 85%). We also see the effect of this process in the distribution of inter-class image similarities between ImageNot and ImageNet (Figure 7 and Appendix B.1 Figure 10); the similarity in images for an ImageNet class tends to be higher than those in an ImageNot class.
Finally, we find that a classifier distinguishes ImageNet and ImageNot well—a common strategy to examine the extent to which two datasets differ (torralba2011unbiased; liu2024decade). We use a ConvNext (liu2022convnet), randomly initialized, as a binary classifier of the two datasets and achieve a 97.0% accuracy when trained on the full training sets. Even with only 1000 samples, accuracy is significantly above chance, and classification accuracy increases with more samples, as shown in Figure 7. Trials include new data resampling and model initialization. We sample training data uniformly for each class except for N/class=1000. The test set is sampled similarly; however, the test images are sampled from the ImageNot/ImageNet test sets.
@@
Similarities in ImageNot and ImageNet WordNet subtrees. Despite the differences in ImageNot/ImageNet, the subset of the WordNet trees spanned by their synsets is structurally similar. They share approximately the same number of nodes and edges, have similar depth, density and modularity, and share a 23% overlap in the parents of the datasets’ synsets – illustrated in Appendix B.1 Figure 12 (more on WordNet subtrees can be found in Appendix B.1 Table 3).
@@ 4 ImageNot preserves model rankings
Our experimental design is centered around evaluating whether ImageNot maintains the model rankings established by ImageNet. Specifically, we want to know: if a new architectural advancement had been evaluated on ImageNot instead of ImageNet, would it still have been recognized as a significant development? Our focus is on a collection of model architectures that represent significant milestones in the evolution of ImageNet models. This includes AlexNet (krizhevsky2012imagenet), VGG (simonyan2014very), Inception V3 (szegedy2015going), ResNet (he2016deep), DenseNet (huang2017densely), MobileNet (howard2019searching), EfficientNet (tan2019efficientnet), Vision Transformers (dosovitskiy2020image), and ConvNeXt (liu2022convnet)—see section 2 for motivating this set. For each model, we select the configuration that achieved the highest reported performance on ImageNet, such as ResNet152 over ResNet18. These models are then trained and evaluated on ImageNot, allowing us to directly compare their rankings with those obtained on ImageNet.
@@
We also investigate ImageNet and ImageNot model performance for finetuning and transfer to the CIFAR10 dataset, which consists of 60,000 32x32 colored images of 10 classes, with 5000 training images per class and 1000 test images per class (cifar-10; krizhevsky2009learning). The learning task is 10-class image classification – Airplane, Automobile, Bird, Cat, Deer, Dog, Frog, Horse, Ship, and Truck. We note that ImageNet contains 50 test images per class (deng2009imagenet) while ImageNot contains 100.
We consider three types of experiments: (i) Random Initialization (From Scratch). We initialize each model with random weights, then train each model on a dataset, and finally, we evaluate each model on the same dataset’s test set. (ii) Fine-Tuning. We initialize each model with pre-trained on ImageNot/ImageNet weights, then we fine-tune the model end-to-end (all weights updated) on CIFAR10, and finally, we evaluate it on CIFAR10’s test set. (iii) Transfer Learning. We initialize each model with pre-trained ImageNot/ImageNet weights, and then we update only the model’s last fully connected layer while training on CIFAR10. Finally, we evaluate it on CIFAR10’s test set.
@@ 4.1 Model development
ImageNet. Our reference rankings are derived from pre-trained models available on Torch Hub (marcel2010torchvision; paszke2019pytorch). These models establish a ranking, which we consider our factual observations. In decreasing order of performance, the ranking is as follows: EfficientNet V2 Large, ConvNeXt Large, Vision Transformer B-16, ResNet 152, Inception V3, DenseNet 161, MobileNet V3 Large, VGG 19, and AlexNet—Figure 1. Please refer to Appendix A.1 for additional specifics regarding these pre-trained models. We also utilize the implementation and pre-trained model weights from torch hub (marcel2010torchvision; paszke2019pytorch) for fine-tuning and transfer with ImageNet pretraining.
On avoiding self-fulfilling prophecies. If not careful, one may fall into a subtle pitfall – the inadvertent confounding effect of knowing the desired ImageNet ranking. Specifically, an experimenter aware of the expected ranking might unconsciously adjust the process until a particular model is in the right place. This iterative tuning could falsely lead to a preservation of the known rankings. To mitigate this risk, we implemented a priori restrictions on the experimenter’s approach for each ranking modality (see Appendix A.1): (1) there is a fixed number of hyperparameter runs that can be launched per model (budget) and (2) the hyper-parameter search space is expanded for all models is expanded only if some model(s) do not train. For ImageNot, we start with the hyperparameters used to train reference ImageNet models to maintain objectivity. However, this means that our resulting training configuration for each model may not be optimal (Figure 8(a)-8(b)) relative to the decade-plus of tuning to obtain the best ImageNet models.
@@ 4.2 Model Rankings
ImageNot preserves model rankings and relative improvements. Figure 1 illustrates that the model rankings between ImageNot and ImageNet are remarkably consistent. We compare each model’s performance improvement relative to AlexNet, the model that kick-started the deep learning era on ImageNet in 2012. This comparison reveals a notable parallel in the progression of accuracy improvements across both ImageNet and ImageNot. Not only do the ordinal rankings of models remain unchanged, but the degree of performance enhancement is also consistent. In simpler terms, this suggests that if a new architecture like EfficientNet (tan2019efficientnet) had been introduced in a counterfactual universe focused on ImageNot instead of ImageNet, we would have witnessed a comparable impressive advancement. The external validity of ImageNet improvements follows. The ImageNet vs. ImageNot relative test accuracies have a linear fit with a slope of 1.01 and a Pearson R of 0.99. We note that these slopes should not be interpreted the same way as those in recht2019imagenet because we compare test accuracies of models with different training tasks/data with different difficulties.
@@
(a) Fine-tuned to CIFAR10—test accuracy compared to standard ImageNet test accuracy for each model architecture
@@
(b) Transfer Learning to CIFAR10—test accuracy compared to standard ImageNet test accuracy for each model architecture.
@@
Figure 8: ImageNet vs. ImageNot. Fine-tuning and Transfer Learning with 95% confidence for regression estimates. Robust corresponds to average accuracy across test-time distribution shifts—error bars do not show for sufficiently small intervals.
@@
Our results also highlight the limitations of relying solely on the external validity of accuracy as the primary metric for the effectiveness of a benchmark. ImageNet model accuracies range from 0.57 (AlexNet) to 0.85 (EfficientNet V2 L), while much noisier ImageNot (Figure 2) model accuracies approximately range from 0.4 (AlexNet) to 0.6 (EfficientNet V2 L). One may look at the drop in accuracy from ImageNet to ImageNot and conclude that ImageNet results have no external validity, but this would be a rather narrow and inaccurate conclusion. From the standpoint of benchmarking, the more relevant observation and notion of external validity is that relative progression in ImageNet accuracy is nearly mirrored by progression in ImageNot accuracy, and model rankings are replicated exactly, despite the significant difference between the two datasets.
@@
For CIFAR10, we do three random restarts and compute the standard errors for each model’s test accuracy. We find that rankings are also preserved for CIFAR10 when training from scratch, up to error bars (Figure 9). kornblith2019better also evaluate CIFAR10 and find a similarly strong correlation between ImageNet accuracy and CIFAR10 accuracy. Given that we opted to include architectures that achieve very similar accuracies on the ImageNet test set (such as EfficientNet-V2-L/ConvNeXt-L or ResNet152/DenseNet161), we expect potential training variance to affect comparisons, particularly given the size of CIFAR10. For instance, when training CIFAR10 from randomly initialized weights or ImageNot (noisy), we observe overlapping error bars for a majority of the best models across random restarts. However, the consistency of relative rankings still holds on average.
@@
ImageNot improvements predict improvements in fine-tuning and transfer learning. When fine-tuning pre-trained ImageNot/ImageNet models on CIFAR10, we observe that improvements in ImageNot predict improvements in finetunability of ImageNot models—similar to ImageNet and ImageNet models (Figure 8(a)). The ImageNot test accuracy and test accuracy of ImageNot models fine-tuned to CIFAR10 have a slope of 0.26 and a Pearson R of 0.83. Finetuned ImageNet has a slope and Pearson R of 0.29 and 0.95, respectively. Here, we find that, as a benchmark (our scope), ImageNot is counterfactually near-equivalent to ImageNet, but as a pretraining dataset, ImageNet is superior.
When transferring ImageNot vs. ImageNet features, we find that ImageNet features are notably more performant (Figure 8(b)). This suggests that the noisiness of ImageNot does have some negative downstream effects on transfer learning. However, we still observe a positive trend in ImageNot test accuracy and transferability of ImageNot models to CIFAR10—slopes of 0.30 and 0.32 with ImageNot and ImageNet, respectively.
Findings hold under robustness. We find that relative improvement also remains closely aligned between ImageNet and ImageNot when we look at average performance across 17 random data augmentations applied to the ImageNot/ImageNet test set (Figure 9; Appendix A.3), e.g., random blurring and cropping. ConvNeXt outperforms EfficientNet V2, however, this holds for both ImageNet and ImageNot, and MobileNet V3 performs marginally better than expected on ImageNot under these augmentations. Nevertheless, our results support the claim that preserving absolute metrics across datasets can be a misleading notion of external validity.
Safety and ethical limitations of web-crawled datasets. While our results suggest that web-crawled datasets with minimal human intervention may be sufficient for benchmarking progress in architecture development, other ethical considerations are a concern. As illustrated by thiel2023identifying, the internet contains many illegal images and images of illegal activity. Particularly, thiel2023identifying previously found evidence of Child Sexual Abuse Material (CSAM) in LAION-5B. We guard against this by using an NSFW language classifier (song2021large) to remove synsets that may be associated with unethical images on top of the default LAION-5B NSFW filter. We did not manually inspect every image in ImageNot, but we employed our model-based filtering and design choices to avoid problematic images as best as possible. We provide ImageNot class names, a detailed description of our generation procedure, and our code so that our results can be easily reproduced.
@@ 5 Conclusion
Despite the drastic differences in ImageNet and ImageNot, we find consistent preservation of model rankings across both datasets. Moreover, this consistency extends to pre-trained models’ fine-tunability and the transferability of pre-trained image embeddings. For benchmarking, we argue that providing robust model rankings, rather than absolute accuracy numbers, should be the primary function of a good benchmark. Consequently, our results suggest that concerns about the external validity of model developments on ImageNet as a benchmark may be overstated.
Many domain experts attribute the rapid advancement in object recognition models to the ImageNet benchmark. From this work, two takeaways emerge to guide and motivate future benchmarks. First, a benchmark that produces robust model rankings may be effective even if accuracy degrades under distribution shifts. Second, clean labels may not be necessary to measure progress effectively.
@@ Appendix A Architectures and Training
AlexNet (krizhevsky2012imagenet) demonstrated the efficacy of deep learning and many architectural advances such as convolutional filters, rectified linear units (ReLUs) as activation functions, and more. Additionally, its relative success on the ILSVRC 2012 is considered to have revitalized interest in neural networks, sparking a wave of research in deep learning. VGG (simonyan2014very) then implemented a simpler and deeper architecture with smaller convolutional filters, along with regularization techniques such as dropout, which yielded improved accuracy and generalization. Inception V3 szegedy2015going focused on multi-scale feature extraction by learning convolutions of different sizes in parallel. Additionally, auxiliary classifiers were added at intermediate layers to enhance gradient flow and act as implicit regularizers. ResNet (he2016deep) further improved upon AlexNet and VGG with the introduction of residual blocks, which allowed for the training of deeper networks while mitigating the vanishing gradient problem and achieving superior performance to its predecessor. DenseNet (huang2017densely) proposed to connect each layer to every other layer to maximize information flow between layers. MobileNet V3 Large focused on optimizing deep learning for mobile and embedded devices using depth-wise separable convolutions and squeeze-and-excitation (SE) modules (hu2018squeeze). EfficientNet (tan2019efficientnet) then improved upon its predecessors with compound and systematic scaling of depth, width, and resolution, as well as combinations of novel optimization and learning strategies, without increasing computational cost. Vision Transformers (ViT-B-16) apply a fundamentally different approach, replacing convolutional feature extractors with self-attention mechanisms; images are treated as sequences of patches, but self-attention across patches allows for long-range dependencies and flexible learning of spatial relationships. ConvNeXt (liu2022convnet) aimed to compete with the state-of-the-art performance of Vision Transformers (ViTs; (dosovitskiy2020image)) by combining components of preceding ConvNet architectures, e.g., residual blocks and kernel sizes.
A.1 ImageNot Models
For ImageNot, we utilize hyperparameters reported for the corresponding ImageNet model to match the model configuration. More details of the hyperparameters used are specified in our Code, with reference to https://github.com/pytorch/vision/tree/main/references/classification. We only change these parameters if models do not train.
AlexNet.
The AlexNet model uses a step scheduler and the SGD optimizer. The learning rate is set to 0.01 with a minimum learning rate of 0.0. The learning rate decay rate is 0.1, with no warmup epochs and an auxiliary learning rate decay after 30 epochs. The model is trained for 90 epochs with a momentum of 0.9 and a weight decay of 0.0001. The batch size is 32.
ConvNeXt Large.
The ConvNeXt Large model uses a cosine annealing with linear warmup scheduler and the AdamW optimizer. The learning rate is 0.001 with a minimum learning rate of 0.0. The learning rate decay rate is 0.1, and warmup is applied for 5 epochs. The model is trained for 600 epochs with a momentum of 0.9 and a weight decay of 0.05. The batch size is 128, and label smoothing is set to 0.1. Additionally, mixup is set to 0.2, cutmix to 1.0, repeated augmentation sampling (RA sampler) is enabled with 4 repetitions, and exponential moving average (EMA) decay is set to 0.99998 with 32 EMA steps.
DenseNet-161.
The DenseNet-161 model uses a step scheduler and the SGD optimizer. The learning rate is 0.1 with a minimum learning rate of 0.0. The learning rate decay rate is 0.01, with no warmup epochs and an auxiliary learning rate decay after 30 epochs. The model is trained for 90 epochs with a momentum of 0.9 and a weight decay of 0.0001. The batch size is 32.
EfficientNet V2-L.
The EfficientNet V2-L model uses a cosine annealing with linear warmup scheduler and the SGD optimizer. The learning rate is 0.5 with a minimum learning rate of 0.0. The learning rate decay rate is 0.1, and warmup is applied for 5 epochs. The model is trained for 600 epochs with a momentum of 0.9 and a weight decay of 0.00002. The batch size is 32, and label smoothing is set to 0.1. Additionally, mixup is set to 0.2, cutmix to 1.0, RA sampler is enabled with 4 repetitions, and EMA decay is set to 0.99998 with 32 EMA steps.
Inception V3.
The Inception V3 model uses a step scheduler and the SGD optimizer. The learning rate is 0.1 with a minimum learning rate of 0.0. The learning rate decay rate is 0.1, with no warmup epochs and an auxiliary learning rate decay after 30 epochs. The model is trained for 90 epochs with a momentum of 0.9 and a weight decay of 0.0001. The batch size is 32. Additionally, the auxiliary loss weight is set to 0.4.
MobileNet V3-Large.
The MobileNet V3-Large model uses a step scheduler and the RMSProp optimizer. The learning rate is 0.064 with a minimum learning rate of 0.0. The learning rate decay rate is 0.973, with no warmup epochs and an auxiliary learning rate decay after 2 epochs. The model is trained for 600 epochs with a momentum of 0.9 and a weight decay of 0.00001. The batch size is 128.
ResNet-152.
The ResNet-152 model uses a step scheduler and the SGD optimizer. The learning rate is 0.1 with a minimum learning rate of 0.0. The learning rate decay rate is 0.1, with no warmup epochs and an auxiliary learning rate decay after 30 epochs. The model is trained for 90 epochs with a momentum of 0.9 and a weight decay of 0.0001. The batch size is 32.
ViT-B/16.
The ViT-B/16 model uses a cosine annealing with a linear warmup scheduler and the AdamW optimizer. The learning rate is 0.001 with a minimum learning rate of 0.0. The learning rate decay rate is 0.033, and warmup is applied for 30 epochs. The model is trained for 300 epochs with a momentum of 0.9 and a weight decay of 0.3. The batch size is 512, and label smoothing is set to 0.11. Additionally, gradient clipping is set to 1.0, mixup is set to 0.2, cutmix to 1.0, RA sampler is enabled with 3 repetitions, and EMA decay is set to 0.99998 with 32 EMA steps.
VGG-19.
The VGG-19 model uses a step scheduler and the SGD optimizer. The learning rate is 0.01 with a minimum learning rate of 0.0. The learning rate decay rate is 0.1, with no warmup epochs and an auxiliary learning rate decay after 30 epochs. The model is trained for 90 epochs with a momentum of 0.9 and a weight decay of 0.0001. The batch size is 32.
We utilize early stopping with patience of 10 based on held-out validation sample accuracy and set a maximum epoch of 1000 for all training. Additionally, to optimize training efficiency and reduce computational costs, we employ the Asynchronous Hyper-Band Scheduler from Ray Tune (li2020system; liaw2018tune), which enables us to further early-stop less promising hyperparameters.
Compute Resource Estimate.
We give a table of approximate times to reproduce our results below.
Table 1: Compute Time. Note that this only accounts for the time it took to train the models used to produce the results in this paper.
Dataset Results GPU-hrs GPUs
ImageNot From Scratch 800 8 A100
CIFAR10 From Scratch 40 2 A100
CIFAR10/ImageNet Finetune 8 2 A100
CIFAR10/ImageNot Finetune 12 2 A100
CIFAR10/ImageNet Transfer Learning 4 2 A100
CIFAR10/ImageNot Transfer Learning 6 2 A100
A.2 CIFAR10 Models